# Sliding DFT-based Signal Recovery for Modulo ADC with 1-bit Folding Information

### v2

@@ II-C Signal Recovery Scheme
Due to the finite bandwidth assumption on $f(t)$, $\hat{F}_{w}(e^{j\frac{2\pi kn}{N}})$ for $\frac{2\pi k}{N}\in(\rho\pi+\Omega_{\mathrm{SL}},2\pi-\rho\pi-\Omega_{\mathrm{SL}})$ can be written as
@@
where the spectral leakage width, denoted $\Omega_{\mathrm{SL}}$, accounts for the frequencies where spectral leakage is still significant. Some energy of the $\underline{\tilde{f}_{w}^{(i)}}[n]$ may leak in the band $(\rho\pi,2\pi-\rho\pi)$ due to the inherent spectral leakage in the DFT computation and cannot be neglected. However, spectral leakage can be controlled by properly choosing the window length $N$ and roll-off parameter $\alpha$. The signals $\underline{z^{(i)}}[n]$ and $\underline{\epsilon^{(i)}}[n]$ are the first-order difference of the modulo residue samples and quantization noise, respectively, in $\tilde{f}_{w}^{(i)}[n]$. The set $\mathcal{S}_{i}$ contains the indices for which $c[n]=1$ in the $i$-th segment. The set identifies the indices in the $i$-th segment for which the DVG changes its output voltage. The samples $z^{(i)}_{\mathcal{S}_{i}}[n]$ (resp. $z^{(i)}_{\mathcal{S}_{i}^{c}}[n]$) corresponds to the value of the modulo residue for all $n$ such that $c[n]=1$ (resp. $c[n]=0$) in the $i$-th segment. The subscript $\mathrm{OOB}$ in $\hat{F}_{\mathrm{OOB}}$ is used to indicate that this is the set of DFT coefficients outside the desired signal bandwidth plus non-negligible spectral leakage.
We are primarily interested in estimating $f^{(i)}[n]$. Our approach is to estimate $z^{(i)}[n]$ first and then subtract $z^{(i)}[n]$ from $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$. This procedure essentially unfolds the modulo ADC output. Let $K$ be the number of discrete frequencies $k$ such that $\frac{2\pi k}{N}\in(\rho\pi+\Omega_{\mathrm{SL}},2\pi-\rho\pi-\Omega_{\mathrm{SL}})$. Equation (II-C) can be written in matrix form as
@@
where $\underline{\mathbf{z}_{w}}\in\mathbb{R}^{N\times 1}$, $\underline{\boldsymbol{\epsilon}_{w}}\in\mathbb{R}^{N\times 1}$, and $\hat{\mathbf{F}}_{\mathrm{OOB}}\in\mathbb{C}^{K\times 1}$ are vectorized form of the finite-length signals $w[n]\cdot\underline{z^{(i)}}[n]$, $w[n]\cdot\underline{\epsilon^{(i)}}[n]$, and $\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})$, respectively. The entry of matrix $\mathbf{V}\in\mathbb{C}^{K\times N}$ in the $k^{\prime}$-th row and $n$-th column is $e^{-j\frac{2\pi nk^{\prime}}{N}}$, where $\frac{2\pi k^{\prime}}{N}$ is the $k^{\prime}$-th discrete frequency in $(\rho\pi+\Omega_{\mathrm{SL}},2\pi-\rho\pi-\Omega_{\mathrm{SL}})$. Let $\underline{\mathbf{z}_{w,\mathcal{S}_{i}}}\in\mathbb{R}^{|\mathcal{S}_{i}|\times 1}$ and $\underline{\boldsymbol{\epsilon}_{w,\mathcal{S}_{i}}}\in\mathbb{R}^{|\mathcal{S}_{i}|\times 1}$ contain the elements of $\underline{\mathbf{z}_{w}}$ and $\underline{\boldsymbol{\epsilon}_{w}}$ whose indices $n$ are in $\mathcal{S}_{i}$. Then, a pre-estimate of $\underline{\mathbf{z}_{w,\mathcal{S}_{i}}}$, denoted as $\underline{\tilde{\mathbf{z}}_{w,\mathcal{S}_{i}}}$, can be expressed as
@@
[…]
Since $z^{(i)}[n]\in 2\lambda^{\prime}\mathbb{Z}$, $\underline{z^{(i)}}[n]\in 2\lambda^{\prime}\mathbb{Z}$. The samples of the first-order difference of modulo residue pre-estimates $\underline{\tilde{z}^{(i)}}[n]$ are rounded to the nearest integer multiple of $2\lambda^{\prime}$ to get the first-order difference of the modulo residue estimates, denoted as $\underline{\hat{z}^{(i)}}[n]$. To undo the first-order difference operation in the modulo residue signal, the formula
@@
is applied. The modulo residue signal estimate $\hat{z}^{(i)}[n]$ is then subtracted from $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$ to unfold the $i$-th segment for $n\in\{0,\cdots,N(1-\frac{\alpha N}{2})\}$. The recovery of the unfolded sequence continues until all segments of $f_{\lambda^{\prime},\mathrm{q}}[n]$ have been processed. If $c[n]=0$ for all $n\in\{0,\cdots,N-1\}$ in the $i$-th segment, we can skip the processing of $i$-th segment and proceed to the next one. Finally, a digital lowpass filter with passband region $\left(-\frac{\Omega_{m}}{2}-\Omega_{\mathrm{SL}},+\frac{\Omega_{m}}{2}+\Omega_{\mathrm{SL}}\right)$ is applied to the result, where $\Omega_{m}=\omega_{m}T_{\mathrm{s}}$. The unfolded signal per segment can be expressed as
@@
$=$ $f^{(i)}[n]+\left\{z^{(i)}[n]-\hat{z}^{(i)}[n]\right\}_{\mathrm{LPF}}+\epsilon^{(i)}_{\mathrm{LPF}}[n]$ (10)
@@ III-A Measuring Recovery Performance
where $N_{0}$ is the signal length. The second line comes from partitioning the original signal into $I$ length-$N$ segments. The third line follows from the expression of the unfolded signal in equation (II-C) and from doing some algebraic manipulation. The fourth line comes from the fact that the quantization noise has zero mean. Thus, the second term in the third line can be removed.
@@ III-B Statistical Model for Quantization Noise
Since MSE involves the mean square of the quantization noise, it is crucial to understand the statistical properties of the quantization noise. The statistical nature of the quantization noise comes from the dithered quantization framework established in [49]. This framework is appropriate for the modulo ADC since the folding operation ensures that the signal does not overload the quantizer. The ‘no overloading’ property of the modulo ADC, together with the triangle dither $d[n]$, guarantees that the quantization noise is a white process and its power does not depend on the input, i.e. $\mathbb{E}\left\{|\epsilon[n]|^{2}\big{|}f_{\lambda^{\prime}}[n]\right\}=\mathbb{E}\left\{|\epsilon[n]|^{2}\right\}$ (see [49, Theorem 2]). For triangle dither $d[n]$ described in Section II, the induced quantization noise has power $\mathbb{E}\left\{|\epsilon[n]|^{2}\right\}=\frac{1}{4}\left(\frac{2\lambda}{2^{b}}\right)^{2}=\frac{\lambda^{2}}{2^{2b}}$.
@@ III-C Incorporating Spectral Leakage
Since spectral leakage cannot be neglected for small $N$, we account for the impact of spectral leakage width $\Omega_{\mathrm{SL}}$ in the proposed recovery algorithm. To this end, we define a quantity $K_{\mathrm{SL}}=\frac{\Omega_{\mathrm{SL}}\cdot N}{\pi}$ called the spectral leakage bin count that corresponds to the number of discrete frequencies in $\left(\rho\pi,\rho\pi+\Omega_{\mathrm{SL}}\right)\cup\left(2\pi-\rho\pi-\Omega_{\mathrm{SL}},2\pi-\rho\pi\right)$. This spectral leakage bin count will appear in the derived MSE performance guarantee. Our goal is to identify how our derived performance guarantee maybe affected by this spectral leakage bin count.
@@ III-D Main Result
For $N$ approaching infinity, equation (13) simplifies to $\mathrm{OF}\geq 3$, which was derived in [14]. The theoretical MSE guarantee established in [14] did not take into account the spectral leakage. As demonstrated in Proposition 2, the spectral leakage bin count $K_{\mathrm{SL}}$ increases the required $\mathrm{OF}$ to satisfy condition (13). Nonetheless, its effect on the OF sufficient condition diminishes fast as $N$ grows.
@@
[…]
$\mathrm{MSE}=\frac{\|f(t)\|_{\infty}^{2}\left(1+\frac{K_{\mathrm{SL}}}{N}\cdot\mathrm{OF}\right)}{\mathrm{OF}\cdot(2^{b}-2)^{2}\cdot\left(\mathrm{OF}\left[1-\frac{K_{\mathrm{SL}}}{N}\right]-2\right)^{2}}.$ (17)
@@
Our result points out an error in [14, Theorem 1]. The derivation of [14, Theorem 1] did not take into account the contribution of the $\epsilon^{(i)}[n]$ for $n\in\mathcal{S}_{i}^{c}$, resulting in $b>3$ condition. In our Theorem 1, the second term in equation (16) accounts for the impact of $\epsilon_{\mathcal{S}_{i}^{c}}^{(i)}[n]$ on the recovery performance. Although we are interested in the modulo residue samples located at $\mathcal{S}_{i}$, $\epsilon_{\mathcal{S}_{i}^{c}}^{(i)}[n]$ still contributes to $\hat{z}^{(i)}[n]$ through the computation of $\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})$. In Section IV, we investigate how the second term of (16) is affected by parameters such as segment length, cardinality of $\mathcal{S}_{i}$, and oversampling factor.
It is also important to point out the detrimental impact of spectral leakage bin count to the MSE performance. Neglecting spectral leakage, equation (17) simplifies to
@@
${\color[rgb]{0,0,0}=\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)}$ (19)
This is similar to the asymptotic growth rate established in [14] which neglected spectral leakage. When $K_{\mathrm{SL}}>0$, the asymptotic growth rate of the MSE becomes $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ only. The $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ growth rate of equation (17) can be maintained if we set $N\geq\mathrm{OF}$. However, using a larger window length increases the computational complexity of the sliding window DFT (see Section III-G).
@@ III-E Comparison with Recovery Guarantees of Existing Methods
We now compare the derived theoretical results with other performance guarantees derived for modulo sampling. One popular recovery procedure is the higher-order differences (HoD) approach developed in [7]. Performance guarantee for this algorithm under a bounded noise setting (e.g., quantization noise) has been analyized in [7, Theorem 6]. More precisely, they showed that for finite-energy bandlimited signals, the noisy unfolded samples can be recovered from noisy modulo samples up to an unknown additive constant, i.e.,
@@
where $p\in\mathbb{Z}$ is not known, if the sampling rate is at least $2^{\alpha}\pi e\times f_{\mathrm{Nyq}}$. Here, $\alpha\in\mathbb{N}$ is a parameter of the bounded noise and depends on both the maximum amplitude of the noise and the modulo threshold. In contrast, Theorem 1 only requires $\mathrm{OF}>3$ under negligible spectral leakage. We also note that $\mathrm{OF}>3$ closely resembles the OF requirement derived in [50] to uniquely identify periodic bandlimited signals under the modulo-DFT sensing model. However, their model first applies DFT on the sampled signal prior to modulo operation. Moreover, the $\mathrm{OF}>3$ result is reminiscent of the spatial oversampling requirement established for DoA estimation using MIMO arrays with modulo nonlinearities [33].
@@
Finally, there exists a line of work [28, 29, 31] which uses time-domain thresholding to unfold the output produced by generalized modulo sampling models. Unlike the DFT-based recovery methods, thresholding is done in the time domain so spectral leakage is not an issue. Moreover, thresholding-based techniques have established recovery guarantees [28, 29]. However, the sequential estimation of the folding parameters requires that the folding times are well-separated in the time domain. Such separation cannot be guaranteed under ideal modulo sampling with zero hysterisis as demonstrated in [29, Lemma 1]. Moreover, the MSE bound established in [29, Proposition 1] is inversely proportional to the hysterisis $h$. Under ideal modulo nonlinearities, this $h$ goes to zero and so the MSE under the thresholding-based recovery becomes unbounded.

### v3

@@ II-C Signal Recovery Scheme
Due to the finite bandwidth assumption on $f(t)$, $\hat{F}_{w}(e^{j\frac{2\pi kn}{N}})$ for $\frac{2\pi k}{N}\in(\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}})$ can be written as
@@
where the spectral leakage width, denoted $\delta_{\mathrm{SL}}$, accounts for the frequencies where spectral leakage is still significant. Some energy of the $\underline{\tilde{f}_{w}^{(i)}}[n]$ may leak in the band $(\rho\pi,2\pi-\rho\pi)$ due to the inherent spectral leakage in the DFT computation and cannot be neglected. Spectral leakage vanishes towards zero for high frequencies and cannot be made zero for any signal. However, spectral leakage can be controlled by properly choosing the window length $N$ and roll-off parameter $\alpha$. The signals $\underline{z^{(i)}}[n]$ and $\underline{\epsilon^{(i)}}[n]$ are the first-order difference of the modulo residue samples and quantization noise, respectively, in $\tilde{f}_{w}^{(i)}[n]$. The set $\mathcal{S}_{i}$ contains the indices for which $c[n]=1$ in the $i$-th segment. The set identifies the indices in the $i$-th segment for which the DVG changes its output voltage. The samples $z^{(i)}_{\mathcal{S}_{i}}[n]$ (resp. $z^{(i)}_{\mathcal{S}_{i}^{c}}[n]$) corresponds to the value of the modulo residue for all $n$ such that $c[n]=1$ (resp. $c[n]=0$) in the $i$-th segment. The subscript $\mathrm{OOB}$ in $\hat{F}_{\mathrm{OOB}}$ is used to indicate that this is the set of DFT coefficients outside the desired signal bandwidth plus non-negligible spectral leakage.
We are primarily interested in estimating $f^{(i)}[n]$. Our approach is to estimate $z^{(i)}[n]$ first and then subtract $z^{(i)}[n]$ from $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$. This procedure essentially unfolds the modulo ADC output. Let $K$ be the number of discrete frequencies $k$ such that $\frac{2\pi k}{N}\in(\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}})$. Equation (II-C) can be written in matrix form as
@@
where $\underline{\mathbf{z}_{w}}\in\mathbb{R}^{N\times 1}$, $\underline{\boldsymbol{\epsilon}_{w}}\in\mathbb{R}^{N\times 1}$, and $\hat{\mathbf{F}}_{\mathrm{OOB}}\in\mathbb{C}^{K\times 1}$ are vectorized form of the finite-length signals $w[n]\cdot\underline{z^{(i)}}[n]$, $w[n]\cdot\underline{\epsilon^{(i)}}[n]$, and $\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})$, respectively. The entry of matrix $\mathbf{V}\in\mathbb{C}^{K\times N}$ in the $k^{\prime}$-th row and $n$-th column is $\frac{1}{\sqrt{N}}e^{-j\frac{2\pi nk}{N}}$, where $\frac{2\pi k}{N}$ is the $k^{\prime}$-th discrete frequency in $(\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}})$. Let $\underline{\mathbf{z}_{w,\mathcal{S}_{i}}}\in\mathbb{R}^{|\mathcal{S}_{i}|\times 1}$ and $\underline{\boldsymbol{\epsilon}_{w,\mathcal{S}_{i}}}\in\mathbb{R}^{|\mathcal{S}_{i}|\times 1}$ contain the elements of $\underline{\mathbf{z}_{w}}$ and $\underline{\boldsymbol{\epsilon}_{w}}$ whose indices $n$ are in $\mathcal{S}_{i}$. Then, a pre-estimate of $\underline{\mathbf{z}_{w,\mathcal{S}_{i}}}$, denoted as $\underline{\tilde{\mathbf{z}}_{w,\mathcal{S}_{i}}}$, can be expressed as
@@
$\underline{\tilde{z}^{(i)}}[n]=\begin{cases}\underline{\tilde{z}^{(i)}_{w}}[n]+\underline{\tilde{z}^{(i-1)}_{w}}\left[n+N\left(1-\frac{\alpha}{2}\right)\right],\;\;0\leq n<\frac{\alpha N}{2}\\ \qquad\qquad\;\;\underline{\tilde{z}^{(i)}_{w}}[n]\;\quad\quad\;,\;\;\frac{\alpha N}{2}\leq n<N\left(1-\frac{\alpha}{2}\right)\end{cases}$
The scaling correction is illustrated in Figure 4. Since $z^{(i)}[n]\in 2\lambda^{\prime}\mathbb{Z}$, $\underline{z^{(i)}}[n]\in 2\lambda^{\prime}\mathbb{Z}$. The samples of the first-order difference of modulo residue pre-estimates $\underline{\tilde{z}^{(i)}}[n]$ are rounded to the nearest integer multiple of $2\lambda^{\prime}$ to get the first-order difference of the modulo residue estimates, denoted as $\underline{\hat{z}^{(i)}}[n]$. To undo the first-order difference operation in the modulo residue signal, the formula
@@
is applied. The modulo residue signal estimate $\hat{z}^{(i)}[n]$ is then subtracted from $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$ to unfold the $i$-th segment for $n\in\{0,\cdots,N(1-\frac{\alpha}{2})\}$. The recovery of the unfolded sequence continues until all segments of $f_{\lambda^{\prime},\mathrm{q}}[n]$ have been processed. If $c[n]=0$ for all $n\in\{0,\cdots,N-1\}$ in the $i$-th segment, we can skip the processing of $i$-th segment and proceed to the next one. Finally, a digital lowpass filter with passband region $\left(-\frac{\Omega_{m}}{2}-\delta_{\mathrm{SL}},+\frac{\Omega_{m}}{2}+\delta_{\mathrm{SL}}\right)$ is applied to the result, where $\Omega_{m}=\omega_{m}T_{\mathrm{s}}$. The unfolded signal per segment can be expressed as
@@
$=$ $f^{(i)}[n]+\left\{z^{(i)}[n]-\hat{z}^{(i)}[n]\right\}_{\mathrm{LPF}}+\epsilon^{(i)}_{\mathrm{LPF}}[n].$ (10)
Figure 4: Illustration of the scaling correction for $\underline{\tilde{z}_{w}^{(i)}}[n]$. The top plot shows the folded samples of the modulo ADC for the (i1)(i-1)-th and $i$-th segments, denoted $f_{\lambda^{\prime},\mathrm{q}}^{(i-1)}[n]$ (blue) and $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$ (red), respectively. The yellow and purple lines correspond to modulo thresholds. The bottom plot shows the first-order differences of the (windowed) modulo pre-estimates in the (i1)(i-1)-th and $i$-th segments, denoted $\underline{\tilde{z}_{w}^{(i-1)}}[n]$ (blue) and $\underline{\tilde{z}_{w}^{(i)}}[n]$ (red), respectively. Due to the tapered cosine window, the recovered amplitudes of $\underline{\tilde{z}_{w}^{(i-1)}}[n]$ and $\underline{\tilde{z}_{w}^{(i)}}[n]$ in the overlap region (maroon box) are attenuated. However, adding the last $\alpha N/2$ samples of $\underline{\tilde{z}_{w}^{(i-1)}}[n]$ to the first $\alpha N/2$ samples of $\underline{\tilde{z}_{w}^{(i)}}[n]$ would correct the amplitude of $\underline{\tilde{z}_{w}^{(i)}}[n]$ from $n\in\{0,\cdots,\frac{\alpha N}{2}-1\}$.
@@ III-A Measuring Recovery Performance
where $N_{0}$ is the signal length. The second line comes from partitioning the original signal into II length-NN segments. The third line follows from the expression of the unfolded signal in equation (II-C) and from doing some algebraic manipulation. The fourth line comes from the fact that the quantization noise has zero mean. Thus, the second term in the third line can be removed.
@@ III-B Statistical Model for Quantization Noise
Since MSE involves the mean square of the quantization noise, it is crucial to understand the statistical properties of the quantization noise. The statistical nature of the quantization noise comes from the dithered quantization framework established in [52]. This framework is appropriate for the modulo ADC since the folding operation ensures that the signal does not overload the quantizer. The ‘no overloading’ property of the modulo ADC, together with the triangle dither $d[n]$, guarantees that the quantization noise is a white process and its power does not depend on the input, i.e. $\mathbb{E}\left\{|\epsilon[n]|^{2}\big{|}f_{\lambda^{\prime}}[n]\right\}=\mathbb{E}\left\{|\epsilon[n]|^{2}\right\}$ (see [52, Theorem 2]). For triangle dither $d[n]$ described in Section II, the induced quantization noise has power $\mathbb{E}\left\{|\epsilon[n]|^{2}\right\}=\frac{1}{4}\left(\frac{2\lambda}{2^{b}}\right)^{2}=\frac{\lambda^{2}}{2^{2b}}$.
@@ III-C Incorporating Spectral Leakage
Since spectral leakage cannot be completely neglected for small $N$, we explicitly account for the effect of non-negligble spectral leakage in the proposed recovery algorithm. To this end, we define the spectral leakage bin count as $K_{\mathrm{SL}}=2\big{\lceil}\frac{\delta_{\mathrm{SL}}\cdot N}{\pi}\big{\rceil}$, which corresponds to the number of discrete frequency bins within the spectral leakage intervals $(\rho\pi,\rho\pi+\delta_{\mathrm{SL}})\cup(2\pi-\rho\pi-\delta_{\mathrm{SL}},2\pi-\rho\pi)$. While $K_{\mathrm{SL}}$ and $\delta_{\mathrm{SL}}$ are generally difficult to compute exactly in practice due to their dependence on the precise shape of the leakage spectrum, we highlight their role in the MSE performance guarantee to reveal how spectral leakage affects recovery accuracy. Finally, we note that the MSE performance guarantee to be derived does not account for the spectral leakage inside $\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}}]$. Nonetheless, our numerical results in Section IV demonstrate that our theoretical MSE predictions still coincide with the simulated MSE despite neglecting some spectral leakage.
@@ III-D Main Result
For $N$ approaching infinity, equation (13) simplifies to $\mathrm{OF}>3$, which was derived in [14]. The theoretical MSE guarantee established in [14] did not take into account the spectral leakage. As demonstrated in Proposition 2, the spectral leakage bin count $K_{\mathrm{SL}}$ increases the required $\mathrm{OF}$ to satisfy condition (13). Nonetheless, its effect on the OF sufficient condition diminishes fast as $N$ grows.
@@
where $M=\underset{i}{\max}\;\|\mathbf{V}_{S_{i}}^{\dagger}\mathbf{V}_{S_{i}^{c}}\|_{\infty}$. The MSE incurred by the proposed recovery algorithm can be written as
$\mathrm{MSE}=\frac{\|f(t)\|_{\infty}^{2}\left(1+\frac{\delta_{\mathrm{SL}}}{\pi}\cdot\mathrm{OF}\right)}{\mathrm{OF}\cdot(2^{b}-2)^{2}\cdot\left(\mathrm{OF}\left[1-\frac{K_{\mathrm{SL}}}{N}\right]-2\right)^{2}}.$ (17)
@@
The derivation of [14, Theorem 1] did not take into account the contribution of the $\epsilon^{(i)}[n]$ for $n\in\mathcal{S}_{i}^{c}$, resulting in $b>3$ condition. In our Theorem 1, the second term in equation (16) accounts for the impact of $\epsilon_{\mathcal{S}_{i}^{c}}^{(i)}[n]$ on the recovery performance. Although we are interested in the modulo residue samples located at $\mathcal{S}_{i}$, $\epsilon_{\mathcal{S}_{i}^{c}}^{(i)}[n]$ still contributes to $\hat{z}^{(i)}[n]$ through the computation of $\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})$. In Section IV, we investigate how the second term of (16) is affected by parameters such as segment length, cardinality of $\mathcal{S}_{i}$, and oversampling factor.
It is also important to point out the detrimental impact of spectral leakage to the MSE performance. Neglecting spectral leakage, i.e. $\delta_{\mathrm{SL}}=0$, equation (17) simplifies to
@@
$=\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ (19)
This is similar to the asymptotic growth rate established in [14] which neglected spectral leakage. When $\delta_{\mathrm{SL}}>0$, the asymptotic growth rate of the MSE becomes $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{2}}\right)$ only. $\delta_{\mathrm{SL}}$ can be reduced if larger window length is used. However, using a larger window length increases the computational complexity of the sliding window DFT (see Section III-G).
@@ III-E Comparison with Recovery Guarantees of Existing Methods
We now compare the derived theoretical results with other performance guarantees derived for modulo sampling. One popular recovery procedure is the higher-order differences (HoD) approach developed in [7]. Performance guarantee for this algorithm under a bounded noise setting (e.g., quantization noise) has been analyzed in [7, Theorem 6]. More precisely, they showed that for finite-energy bandlimited signals, the noisy unfolded samples can be recovered from noisy modulo samples up to an unknown additive constant, i.e.,
@@
where $p\in\mathbb{Z}$ is not known, if the sampling rate is at least $2^{\alpha}\pi e\times f_{\mathrm{Nyq}}$. Here, $\alpha\in\mathbb{N}$ is a parameter of the bounded noise and depends on both the maximum amplitude of the noise and the modulo threshold. In contrast, Theorem 1 only requires $\mathrm{OF}>3$ under negligible spectral leakage. We also note that $\mathrm{OF}>3$ closely resembles the OF requirement derived in [53] to uniquely identify periodic bandlimited signals under the modulo-DFT sensing model. However, their model first applies DFT on the sampled signal prior to modulo operation. Moreover, the $\mathrm{OF}>3$ result is equivalent to Itoh’s condition [54] and is reminiscent of the spatial oversampling requirement established for DoA estimation using MIMO arrays with modulo nonlinearities [33].
@@
Finally, there exists a line of work [28, 29, 31] which uses time-domain thresholding to unfold the output produced by generalized modulo sampling models. Unlike the DFT-based recovery methods, thresholding is done in the time domain so spectral leakage is not an issue. Moreover, thresholding-based techniques have established recovery guarantees [28, 29]. However, the sequential estimation of the folding parameters requires that the folding times are well-separated in the time domain. Such separation cannot be guaranteed under ideal modulo sampling with zero hysteresis as demonstrated in [29, Lemma 1]. Moreover, the MSE bound established in [29, Proposition 1] is inversely proportional to the hysteresis $h$. Under ideal modulo nonlinearities, this $h$ goes to zero and so the MSE under the thresholding-based recovery becomes unbounded.